# Bidirectional Consistency Models

### v2

@@ 4.2 Inversion and Reconstruction
(a) Unconditional model
(b) Conditional model
Figure 4: Comparison of reconstruction error between DDIM, EDM, and our proposed BCM. Both conditional and unconditional BCM can achieve a low error with significantly fewer NFEs.
As highlighted earlier, a distinctive feature of BCM is its ability to invert an input image ${\mathbf{x}}_{0}$ back to a noise image ${\mathbf{x}}_{T}$, and then reconstruct $\hat{{\mathbf{x}}}_{0}$ from ${\mathbf{x}}_{T}$ using notably few NFEs. To demonstrate this capability, we evaluate the reconstruction performance on the CIFAR-10 test set. A subset of the training set, containing 2,000 samples, was used to adjust the hyperparameters, which we found to generalize effectively to the test set. We report the per-dimension mean squared error (scaled to [0,1]), along with ODE-based baselines, including DDIM [6] and EDM [15]  [^2] For DDIM, we use the reported MSE by Song et al. [6]; for EDM, we use the checkpoint provided in the official implementation at https://github.com/NVlabs/edm?tab=readme-ov-file, and re-implement a deterministic ODE solver following Algorithm 1 in [15]. in Figure 4. To further demonstrate our model’s applicability, we also train a conditional BCM and report its reconstruction performance in Figure 4(b). The results indicate that BCM achieves a lower reconstruction error than ODE-based diffusion models, with significantly fewer NFEs.
We visualize the noise image generated by BCM at Figure 15 in Appendix F, from which we can see that BCM Gaussianizes the input image as desired. We also provide examples of the reconstruction in Figure 16 in Appendix F. We observe that using only 1 NFE for inversion sometimes introduces very slight mosaic artifacts, possibly because both endpoints of the time horizon are less effectively covered and learned during training. Fortunately, the artifact is well suppressed when using more than 2 NFEs. Also, we find that the inversion process can occasionally alter the image context. However, it is important to note that this phenomenon is not unique to our model but also occurs in ODE-based diffusion models, such as EDM (see Figures 16(f) and 16(g) for examples), even though they employ more NFEs.
@@ 4.3 Applications with Bidirectional Consistency
We first remark that since our proposed BCM inherits the properties of CMs, it can also handle the applications introduced by Song et al. [17] similarly, including colorization, super-resolution, stroke-guided image generation, denoising, and interpolation between two generated images. Therefore, our discussion primarily centers on the applications enhanced or made possible by BCM’s bidirectional consistency. Specifically, we showcase that BCM can achieve superior inpainting quality and interpolate two real images. We also demonstrate its potential in blind restoration of compressed images and defending black-box adversarial attack. We provide experimental details in Section A.3.
4.3.1 Interpolation
While CMs can perform interpolations between two generated images [17], our BCM can interpolate between two given real images, which is a more meaningful application. Specifically, we can first invert the given images into noises, smoothly interpolate between them, and then map the noises back to images. We illustrate some examples in Figure 5 and more results in Figure 17.
Here, we note a caveat in the implementation of BCM’s interpolation: recall that when inverting an image, we first inject a small initial noise, as described in Section 3.4. In the context of inpainting, we find it crucial to inject different initial noises for each of the two given images to avoid sub-optimal results, as illustrated in Figure 18. We defer further discussions to Section A.3.1.
4.3.2 Inpainting
Given an image with a masked region, our model can perform inpainting leveraging its bidirectional consistency. Specifically, given an image with some pixels missing, we first add a small noise to the missing pixels. Then, we start to invert the image by multi-step inversion. Different from our standard inversion algorithm in Algorithm 5, at each iteration, we manually replace the region of missing pixels with a new Gaussian noise, whose scale corresponds to the current time step. In this way, we gradually change the out-of-distribution (OOD) region to in-distribution noise. In the end, we map the entire noisy image back to time 00, finishing the inpainting. We summarize this process in Algorithm 6, and illustrate our BCM inpainting results in Figures 6(b) and 19.
Song et al. [17] also demonstrated that CMs could also perform image inpainting in an iterative way. We thus include the inpainting results of CMs for comparison in Figures 6(c) and 20. As we can see, CMs produce worse inpainting compared to our BCM in most cases, which is also reflected by the test set FID. Also, we note that BCM’s inpainting requires much fewer NFEs than theirs. A possible reason for CMs’ inferior outcomes is that in CMs, we can only add noise manually, and hence we face a dilemma during inpainting: on the one hand, we need large noise to fill in the missing region; on the other hand, a large amount of noise (added to the whole image) engulfs the remaining image signals, causing the inpainted content to not align well with the surrounding areas. BCM, however, circumvents these issues by employing bidirectional consistency to achieve nearly deterministic inversion. Therefore, the missing and remaining regions can align well in the noise space, while still mostly preserving the contents.
Figure 5: Interpolation between two real images.
@@ 4.3.2 Inpainting
(b) Inpainting with BCM (NFE=4, test set FID 12.32)
(c) Inpainting with CM (NFE=18, test set FID 13.16)
Figure 6: Comparison between image inpainting by BCM and CMs.
4.3.3 Blind Restoration of Compressed Image
Interestingly, BCM can perform blind restoration of compressed images by inverting an image with compression artifacts, ${\mathbf{x}}_{0}$, back to ${\mathbf{x}}_{t}$, and then reconstructing $\hat{{\mathbf{x}}}_{0}$ from ${\mathbf{x}}_{t}$. During inversion, the model maps images with OOD artifacts to in-distribution noise, effectively eliminating these artifacts. The hyperparameter $t$ allows control over this process: a larger $t$ enhances restoration but risks altering the context.
We evaluate the BCM’s restoration performance on two types of compression artifacts: one resulting from the traditional JPEG codec and the other from a more recent implicit neural representation (INR)-based codec [30]. We visualize the outcomes in Figure 7 and provide more examples in Figures 21 and 22. JPEG compression often leads to discrete cosine transform (DCT) blocks and some color shift, whereas INR compression tends to over-smooth details and may introduce high-frequency noises due to its sine activation function [31]. Surprisingly, we can see that BCM effectively eliminates these different artifacts.
However, we also observe that BCM occasionally alters the content. We can reduce the hyperparameter $t$ for a trade-off. Alternatively, we can also perform several optimization steps to minimize the difference between the reconstructed and compressed images, akin to the strategy proposed by Xu et al. [32] and Mokady et al. [12]. BCM still holds an advantage in this context over previous diffusion models, as each optimization step requires significantly fewer NFEs, enhancing speed and practicality. As our experiments are mainly intended as a proof of concept, we leave this direction for further exploration.
(a) Restoration of JPEG-compressed Images.
(b) Restoration of INR-compressed Images.
Figure 7: Blind restoration of compressed images. The first row is the compressed images, while the second is the restoration results. These results are cherry-picked to highlight our model’s features. We also provide non-curated results in Figures 21 and 22 in Appendix. We can see that while JPEG and INR-based codecs introduce different artifacts, BCM performs well in eliminating both of them. However, we also note that our model sometimes alters the content in an undesired manner.
4.3.4 Defending Black-Box Adversarial Attack
Table 3: Accuracies (%) under black-box adversarial attack among two source models and four unsecured target models on CIFAR-10, before and after purifying with BCM.
Source Model PGD Budget PARN-18 WRN-34-10 VGG-16 MobileNet-V2
before after before after before after before after
PreActResNet-18 $\varepsilon_{\infty}=0/255$ 92.64 70.35 95.01 71.33 93.18 71.45 91.26 69.67
$\varepsilon_{\infty}=4/255$ 34.20 66.77 32.24 68.65 57.50 69.34 35.11 66.68
$\varepsilon_{\infty}=8/255$ 0.00 64.24 8.00 66.85 26.54 67.51 10.77 64.86
$\varepsilon_{\infty}=12/255$ 0.00 61.83 3.31 64.79 13.16 65.69 4.97 62.69
$\varepsilon_{\infty}=16/255$ 0.00 59.83 1.54 62.07 7.79 63.37 3.01 60.88
$\varepsilon_{\infty}=24/255$ 0.00 53.56 0.48 56.49 3.72 56.99 1.57 54.81
$\varepsilon_{\infty}=32/255$ 0.00 44.94 0.24 46.63 2.32 47.71 1.10 45.41
WideResNet-34-10 $\varepsilon_{\infty}=0/255$ 92.64 70.35 95.01 71.33 93.18 71.45 91.26 69.67
$\varepsilon_{\infty}=4/255$ 40.59 68.05 6.77 68.20 64.22 69.30 42.74 67.32
$\varepsilon_{\infty}=8/255$ 11.70 66.51 0.04 66.19 35.62 67.96 15.48 65.38
$\varepsilon_{\infty}=12/255$ 4.90 64.68 0.00 64.35 20.60 66.67 7.78 64.11
$\varepsilon_{\infty}=16/255$ 2.50 62.86 0.00 62.07 12.78 63.82 5.50 61.41
$\varepsilon_{\infty}=24/255$ 1.44 56.59 0.00 55.79 7.22 57.96 3.27 56.12
$\varepsilon_{\infty}=32/255$ 0.88 47.66 0.00 45.91 4.54 49.45 2.59 46.81
Deep neural classifiers are known to be vulnerable to adversarial examples [33, 34]. These malicious examples are created by adding adversarial noises almost imperceptible to human eyes (e.g., within a range of $\|\varepsilon\|_{\infty}\leq 8/255$ for RGB images within $[0,1]$) to clean examples, usually performed by doing gradient ascent on a source model. In the real world, because weights of target models are usually unknown, it becomes a common practice to conduct black-box adversarial attacks, i.e., generating adversarial examples on source model $\mathcal{A}$ and using them to attack an unsecured target model $\mathcal{B}$. These adversarial examples typically transfer well (i.e., remain adversarial) by a large percentage, even across very different network capacities or architectures [35].
In this section, we investigate the idea of using our BCM to defend against such an attack. The main idea is to use our BCM to purify potentially adversarial examples before inputting them into an unsecured classifier. Specifically, given a potentially adversarial image, we first invert it to the noise space of $T=80$ and then map it back to the image space with the help of bidirectional consistency. To verify the effectiveness of such purification, we create adversarial examples on two source models (PreActResNet-18 [36] and WideResNet-34-10 [37]) to attack four unsecured target models (PARN-18 [36], WRN-34-10 [37], VGG-16 [38] and MobileNet-V2 [39]) with $\ell_{\infty}$ norm PGD attack [40]. Importantly, since the models being attacked are blind to the attack budget (e.g., the adversarial noise scale) [^3] We remark that this is the key reason why CMs are not compatible with our black-box adversarial defense setting: without inversion ability, the adversarial noise scale has to be known for proper denoising, which is impossible in most real-world cases., we vary the adversarial noise scale over $\varepsilon_{\infty}\in[0/255,32/255]$ on the CIFAR-10 test set to examine our approach’s performance in defending attacks of variant strengths. We summarize the results in Table 3.
First, while the adversarial examples show impressive transferability across different networks, e.g., accuracies decrease to $<36\%$ when $\varepsilon_{\infty}=8/255$ for all target models, BCM’s purification significantly improves the defense against such black-box adversarial attacks, e.g., improving accuracies to $>64\%$ at $\varepsilon_{\infty}=8/255$. Second, we observe that our method is very robust to variant attack budgets that even under extremely strong attacks of $\varepsilon_{\infty}=32/255$ where all target models show $<5\%$ accuracies without purification, our approach still holds $>44\%$ accuracies for all models.
However, same as the observations made in Section 4.3.3, BCM’s reconstruction (or here, purification) sometimes alters image contents, which is an undesired property for classification. This explains why the accuracies drop to $\sim 70\%$ after reconstruction even for $\varepsilon_{\infty}=0/255$ (clean samples). One feasible remedy is to invert the image to a smaller noise scale, i.e., to ${\mathbf{x}}_{t},t<T$, but at the cost of insufficient defense strength when the attack budget is large. We present this trade-off in Figure 13.
@@ 5 Related Works
Accelerating DMs’ Generation. Before Song et al. [17] introduced consistency models, several attempts were made to accelerate the generation of diffusion models. Existing methods include faster ODE solvers [6, 41, 42] and distillation [43, 44, 27, 17]. However, these ODE solvers generally require over 10 steps for satisfactory samples, and the distillation approach caps the performance at the level of the pre-trained diffusion model, as highlighted by Song and Dhariwal [18].
Inversion of Generation. Inversion of an input image is a key step in downstream tasks like image editing [12, 45, 14, 13]. Current methods typically include encoders [46, 47], optimization [12, 48], and model fine-tuning or modulation [49, 50]. For diffusion-based models, the most common inversion method is based on the inversion ODE solver, as first proposed by Song et al. [6]. However, since the ODE solver relies on local linearization, the inversion typically provides an insufficient reconstruction of the original input, especially when using fewer diffusion steps. Wallace et al. [45] proposed Exact Diffusion Inversion via Coupled Transformations (EDICT), addressing this issue by maintaining two coupled noise vectors and using them to invert each other in an alternating fashion. However, the computational time is approximately doubled. Mokady et al. [12] provide another solution by performing gradient optimization using the ODE inversion outcome as an initialization. Our proposed method, on the other hand, can provide better inversion than the ODE baselines in only a handful of steps, and is also compatible with Mokady et al. [12]’s approach.
@@ 6 Conclusions and Limitations
In this work, we introduce the Bidirectional Consistency Model (BCM), enhancing upon existing consistency models [17, 18] by establishing a stronger consistency. This consistency ensures that points along the same trajectory of the probability flow (PF) ODE map to each other, thereby unifying generation and inversion tasks within one framework. By exploiting its bidirectional consistency, we devise new sampling schemes and showcase applications in a variety of downstream tasks. Given the pervasive application of diffusion models and the comparatively limited research into accelerating their inversion tasks, we believe that our model opens a new avenue for further exploration.
@@
A main limitation of our method is that while employing more steps in generation or inversion can initially enhance results, the performance improvements tend to plateau quickly. Consequently, increasing the Number of Function Evaluations (NFEs) beyond a certain point does not yield further performance gains. This phenomenon could be attributed to the fact that while increasing NFEs reduces the error at each step, cumulatively, more steps can lead to an accumulation of errors. A potential solution involves employing the parameterization and tricks proposed by Kim et al. [19].
Additionally, as previously mentioned, our method delivers imperfect inversion, sometimes altering the image content. While this also happens in ODE-based methods [15, 6], it complicates the downstream applications. Therefore, future work can involve developing more accurate inversion techniques, such as the approach proposed by Wallace et al. [45].
Another limitation of our method is its large training requirements, particularly when processing higher-resolution images. A potential approach to mitigate this issue is distillation, inspired by the consistency distillation introduced by Song et al. [17]. Alternatively, we can apply BCM in the latent space, similar to the latent consistency model proposed by Luo et al. [51].
Appendix A Experiment Details
A.1 Training Settings
We use a batch size of 1,024 with the student EMA decay rate of $0.99993$ for 400k iterations with RAdam optimizer [52] using learning rate $0.0001$. We use the NCSN++ network architecture proposed by Song et al. [7], with the modification described in Section 3.1. Regarding other training settings, including the scheduler function for $N(\cdot)$, the sampling probability for $t_{n}$ (aka the noise schedule $p(n)$ in [17, 18]), the distance measure $d(\cdot,\cdot)$, we follow exactly Song and Dhariwal [18], and restate below for completeness’s sake:
$K$ is the total training iterations
A.2 Sampling and Inversion Configurations
Sampling. For unconditional BCM, we use ancestral sampling with $t_{1}=1.2$ for NFE$=2$, zigzag sampling with $\varepsilon_{1}=0.2,t_{1}=0.8$ for NFE$=3$ and the combination of ancestral sampling and zigzag sampling with $t_{1}=1.2,\varepsilon_{1}=0.1,\tau_{1}=0.3$ for NFE$=4$. For BCM-deep, we use ancestral sampling with $t_{1}=0.7$ for NFE$=2$, zigzag sampling with $\varepsilon_{1}=0.4,t_{1}=0.8$ for NFE$=3$ and the combination with $t_{1}=0.6,\varepsilon_{1}=0.14,\tau_{1}=0.3$ for NFE$=4$.
Inversion. For both unconditional and conditional BCM, we set $\varepsilon=t_{1}=0.07$, $t_{2}=6.0$ and $t_{3}=80.0$ for NFE$=2$ and $t_{2}=1.5,t_{3}=4.0,t_{4}=10.0,t_{5}=80.0$ for NFE$=4$. We use 1-step generation to map the inverted noise to reconstructed images and evaluate the per-dimension MSE between the original images and their reconstructed counterparts.
@@ A.2 Sampling and Inversion Configurations
A.3 Applications with Bidirectional Consistency
Here we provide more details about the applications we demonstrated in Section 4.3.
A.3.1 Interpolation
We first invert the two given images ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ to noise at $T=80.0$ using Algorithm 5. To avoid subscript overloading, in this section, we denote their noise as ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$, respectively. Specifically, we find that adopting a 3-step inversion with $\varepsilon=t_{1}=0.07,t_{2}=1.5,t_{3}=6.0$, and $t_{4}=80.0$ is sufficient for good reconstruction results. Note, that we use two different small initial noise, i.e., $\bm{\sigma}_{1},\bm{\sigma}_{2}\sim\mathcal{N}(0,{\bm{I}}),\bm{\sigma}_{1}\neq\bm{\sigma}_{2}$ when inverting ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ respectively.
Since BCMs learn to amplify the two initial Gaussian i.i.d. noises, it is reasonable to hypothesize that the amplified noises (i.e., the embeddings) ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$ reside on the same hyperspherical surface as if ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$ are directly sampled from $\mathcal{N}(0,T^{2}{\bm{I}})$. Therefore, following Song et al. [17], we use spherical linear interpolation as
As we discussed in the main text, it is crucial to set $\bm{\sigma}_{1}\neq\bm{\sigma}_{2}$ for inversion. A possible reason is that if using $\bm{\sigma}_{1}=\bm{\sigma}_{2}$ for inversion, the inverted noises ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$ may reside on an unknown submanifold instead of the hyperspherical surface of Gaussian, and hence Equation 23 cannot yield ideal interpolation results. In Appendix E, we present visualization and discussions on the geometric properties of the noise space.
A.3.2 Inpainting

### v4

@@ 4.2 Inversion and Reconstruction
We evaluate BCM’s capability for inversion on CIFAR-10 and ImageNet-64, and report the per-dimension mean squared error (scaled to [0,1]) in Figure 7. We also present ODE-based baselines, including DDIM (Song et al., 2021a) and EDM (Karras et al., 2022) for comparison. The hyperparameters for inversion are tuned on a small subset of the training set and are discussed in Section C.2. We can see BCM achieves a lower reconstruction error than ODE-based diffusion models, with significantly fewer NFEs, on both data sets.
We visualize the noise generated by BCM at Figure 16 in Appendix G, from which we can see that BCM Gaussianizes the input image as desired. We also provide examples of the reconstruction in Figure 17 in Appendix G. We observe that using only 1 NFE for inversion sometimes introduces slight mosaic artifacts, possibly because both endpoints of the time horizon are less effectively covered and learned during training. Fortunately, the artifact is well suppressed when using more than 1 NFEs. Also, we find that the inversion process can occasionally alter the image context. However, it is important to note that this is not unique to BCM but also occurs in ODE-based diffusion models, such as EDM (see Figures 17(f) and 17(g) for examples), even though they employ more NFEs.
@@ 4.3 Applications with Bidirectional Consistency
(a) CIFAR-10.
(b) ImageNet-64.
Figure 5: Reconstruction MSE on CIFAR-10 & ImageNet.
Figure 6: Interpolation between two real images. We include more results in figs. 18, 20 and 21.
@@
(b) BCM’s inpainting (NFE=4, FID 12.32). More examples in fig. 22.
(c) CM’s Inpainting (NFE=18, FID 13.16). More examples in fig. 23.
Figure 7: Inpainting by BCM and CMs on CIFAR-10 test set.
Since BCM inherits the consistency of CMs, it can handle the applications introduced by Song et al. (2023) in the same way, including colorization, super-resolution, stroke-guided image generation, denoising, and interpolation between two generated images. However, BCM’s unique bidirectional consistency can enhance or enable more applications. In this section, we showcase that BCM can interpolate two real images and achieve superior inpainting quality with fewer NFEs. We also demonstrate its downstream applications for blind restoration of compressed images.
Interpolation. While CMs can perform interpolations between two generated images (Song et al., 2023), our BCM can interpolate between two given real images, which is a more meaningful application. Specifically, we can first invert the given images into noises, smoothly interpolate between them, and then map the noises back to images. We illustrate some examples in Figure 7. Here, we note a caveat in the implementation of BCM’s interpolation: recall that when inverting an image, we first inject a small initial noise, as described in Section 3.4. In the context of interpolation, we find it crucial to inject different initial noises for each of the two given images to avoid sub-optimal results, as illustrated in Figure 19. We defer further discussions to Section C.3.1.
Inpainting. BCM’s bidirectional consistency can also help with inpainting: having an image with some pixels missing, we first add a small noise to the missing pixels and then invert the image by multi-step inversion. Different from our standard inversion algorithm in Algorithm 5, at each iteration, we manually replace the region of missing pixels with a new Gaussian noise, whose scale corresponds to the current time step. In this way, we gradually fill in the missing region to in-distribution noise. In the end, we map the entire noisy image back to time 00, finishing the inpainting. We summarize this process in Algorithm 6, and include more details on the hyperparameters in Section C.3.2. We illustrate our BCM inpainting results on CIFAR-10 in Figure 7, with the CM’s inpainting results using the algorithms proposed by Song et al. (2023) (Algorithm 4 in (Song et al., 2023)) for comparison. Our method produces better inpainting results within much fewer NFEs.
@@ 5 Related Works
Accelerating Diffusion’s Generation. In addition to CM and CTM we discussed in Section 2, many attempts have also been made to accelerate the generation of DMs. Existing methods include faster ODE solvers (Song et al., 2021a; Zhang & Chen, 2022; Lu et al., 2022b; Zhou et al., 2023) and distillation (Salimans & Ho, 2022; Luhman & Luhman, 2021; Zheng et al., 2023; Luo et al., 2024; Xie et al., 2024). However, these ODE solvers generally require 5-10 steps for satisfactory samples, and the distillation approach may cap the performance at the level of the pretrained diffusion model, as highlighted by Song & Dhariwal (2024). On the other hand, while distillation approaches have achieved impressive generation quality, an intrinsic property of CMs is their ability to enable fast traversal along ODEs. This allows for the design of broader applications; for example, Zhang et al. (2024) leverage this property to develop an efficient importance sampler. Our proposed BCM further enriches this by enabling both forward and backward traversal.
Inversion. Inversion of an input image is a key step in downstream tasks like image editing (Mokady et al., 2023; Wallace et al., 2023; Hertz et al., 2023; Huberman-Spiegelglas et al., 2023). Current methods typically include encoders (Richardson et al., 2021; Tov et al., 2021), optimization (Mokady et al., 2023; Abdal et al., 2019), and model fine-tuning or modulation (Roich et al., 2022; Alaluf et al., 2022). For diffusion-based models, the most common inversion method is based on solving PF ODE backward, as first proposed by Song et al. (2021a). However, since the ODE solver relies on local linearization, this ODE-based inversion typically provides an insufficient reconstruction, especially when using fewer diffusion steps. One way to address this issue is to perform gradient optimization using the ODE-based inversion as an initialization, as introduced by Mokady et al. (2023). Our proposed BCM can provide better inversion with fewer NFEs and is compatible with Mokady et al. (2023)’s approach for a better reconstruction quality.
In a concurrent work, Starodubcev et al. (2024) propose invertible consistency distillation, which trains a second network to learn the inverse process and demonstrate its application in text-guided image editing. While both approaches are motivated by similar goals, BCM differs from theirs in methodology by using a single network to learn the integration between any two points along the PF ODE, effectively integrating both generation and inversion without increasing the model size. Our proposed objective and sampling scheme allow us to achieve bidirectional consistency while maintaining or even improving generation performance.
@@ 6 Conclusions and Limitations
In this work, we introduce the Bidirectional Consistency Model (BCM), enhancing upon existing consistency models (Song et al., 2023; Song & Dhariwal, 2024; Kim et al., 2024) by establishing a stronger consistency. This consistency ensures that points along the same trajectory of the probability flow (PF) ODE map to each other, thereby unifying generation and inversion tasks within one framework. By exploiting its bidirectional consistency, we devise new sampling schemes and showcase applications in a variety of downstream tasks. Unlike other distillation approaches, a notable property of CMs is their ability to enable fast traversal along the entire PF ODE. Our proposed BCM further enriches this, and we believe that it will open a new avenue for further exploration.
@@
One limitation of our method is that while employing more steps in generation or inversion can initially enhance results, the performance improvements tend to plateau quickly. Increasing the Number of Function Evaluations (NFEs) beyond a certain point does not yield further performance gains. This is similar to CMs, where there is no performance gain with more than 2 NFEs. A potential solution involves employing the parameterization and tricks proposed by Kim et al. (2024). Additionally, our method delivers imperfect inversion, which sometimes alters the image content. However, we should note that this also happens in ODE-based DMs. Future work can involve developing more accurate inversion techniques, like the approach by Wallace et al. (2023). Alternatively, we can fine-tune our BCM for specific downstream tasks to further enhance its performance.
Broader Impact and Ethics Statements
BCM can accelerate the generation and inversion of diffusion models, which may pose a risk of generating or editing images with harmful or inappropriate content, such as deepfake images or offensive material. Strong content filtering or even regulatory rules are required to prevent the creation of unethical or harmful content.
Acknowledgment
We are grateful to Zongyu Guo for the helpful discussion throughout this project, especially for pointing out the significance of the inverse process, suggesting some potential applications, and proofreading the preprint version of this work. We also thank Wenlin Chen and Sergio Calvo-Ordoñez for reviewing and providing feedback on the manuscript. JH acknowledges support from Prof. José Miguel Hernández-Lobato’s Turing AI Fellowship under grant EP/V023756/1.
Appendix A Algorithms
@@
Algorithm 4 Combination of ancestral and zigzag sampling
Generated image ${\mathbf{x}}$.
${\mathbf{x}}_{t_{N}}\leftarrow{\mathbf{x}}_{T}$.
# Ancestral sampling steps
for $n=N,\cdots,2$ do
end for
# Zigzag sampling steps
${\mathbf{x}}_{\tau_{M}}\leftarrow{\mathbf{x}}_{t_{1}}$.
for $m=M,\cdots,2$ do
end for
${\mathbf{x}}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{\tau_{1}},\tau_{1},0)$.
Return: ${\mathbf{x}}$.
@@
   (Here, we exemplify with 1-step sampling, but note that multi-step schemes can also be used)